OpenStudy (anonymous):

Each unit on the map represents 5 miles. What is the actual distance from Oceanfront to Seaside?

OpenStudy (anonymous):

can you upload the map?

OpenStudy (anonymous):

can you give me a chance

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

if i draw up a sketch like this, can you figure out the distance of 'x' and 'y' i put on the triangle that I made from the graph? 

OpenStudy (luigi0210):

 We know those sides, now we have to use \(A^2+B^2=C^2\) to find the remaining.

OpenStudy (anonymous):

That's a 6 and a 5 ? correct ? @Luigi0210

OpenStudy (luigi0210):

Yup~

OpenStudy (anonymous):

61=c squared

OpenStudy (luigi0210):

Right, now take the square root of both sides. Giving you \(\pm \sqrt{61}=c\).. but we can't have a negative distance for this so just \(\sqrt{61}=c\) works c:

OpenStudy (anonymous):

it's either 10, 8 or 40 miles

OpenStudy (anonymous):

@texaschic101 @TERM_THE_INATOR @agent0smith

OpenStudy (anonymous):

@Shay17 we all want to help you, but we do need you to do the work, what were you able to calculate from what luigi helped u with?

OpenStudy (anonymous):

you can clearly see the answer i came up with and it was obviously wrong i'm not still on this question for nothing i have no idea what the answer is @DemolisionWolf

OpenStudy (anonymous):

@Jack1 can you help me with this one also ?

OpenStudy (jack1):

so u had it all the way before, u just need to round it to the closest answer use pythag's theory as its a right angled triangle C^2 = A ^2 = B^2 a = 5, b = 6 so c^2 = 36 + 25 = 61 so c = sqrt 61 = 7.8... so with rounding that comes to about 8

OpenStudy (jack1):

remember to give @Luigi0210 a medal for this one tho, he put in the effort with his drawings and explanations above

OpenStudy (jack1):

so does that all make sense now @Shay17 ?

OpenStudy (anonymous):

yes thanks again @Jack1